OpenStudy (anonymous):

Simplify and find all numbers that must be excluded from the domain of the simplified rational expression. (x^2-x)/(x+5) times (x^2 + 5x + 4)/(x^4 - x^2)

OpenStudy (e.mccormick):

x^2-x/x+5 times x^2 + 5x = 4/x^4 - x^2 \(\leftarrow\) needs some ( ) to make the order clear. What is over what?

OpenStudy (anonymous):

Edited, Thank you.

OpenStudy (e.mccormick):

The = is inside the ( )???

OpenStudy (anonymous):

It was meant to be an + ..

OpenStudy (e.mccormick):

AH! OK. That makes sense since they are the same key. \(\dfrac{x^2-x}{x+5} \times \dfrac{x^2 + 5x + 4}{x^4 - x^2}\)

OpenStudy (anonymous):

Correct.

OpenStudy (e.mccormick):

OK. Because this is multiplication of fractions, the easiest thing to do is factor it and cancel. But leave it in factored form for a bit to answer the excluded domain part.